# A HHO formulation for variable density incompressible flows where the density is purely advected

### v1

@@ 4.1 Kovasznay test case
As a first validation step we seek to numerically demonstrate pure advection of the density by takling the Kovasznay flow problem [18]. The analytical solution of the incompressible Navier–Stoke equations, which describes steady flow behing a grid made of equally spaces parallel rods for a fluid with constant density, reads as follows
$\bm{u}$ $=\left[1-\mathrm{e}^{\kappa x}\cos{\left(2\pi y\right)}\right]\bm{i}+\dfrac{\kappa}{2\pi}\mathrm{e}^{\kappa x}\sin{\left(2\pi y\right)}\bm{j},$ (21)
$p$ $=p_{0}-\dfrac{1}{2}\mathrm{e}^{2\kappa x},$
$\rho$ $=1,$
$\kappa=\dfrac{\mathrm{Re}}{2}-\sqrt{\dfrac{\mathrm{Re}^{2}}{4}+4\pi^{2}}.$
The Kovasznay flow problem is defined on the bi-unit square computational domain $\Omega=\left(-0.5,1.5\right){\times}\left(0,2\right)$ and boundary conditions are derived from the Kovasznay analytical solution (21) at $\mathrm{Re}{=}\frac{1}{\mu}{=}40$. In particular, a Neumann boundary condition is imposed at the outflow boundary (right side of the square domain) while Dirichlet boundary conditions are imposed on the remaining boundaries. Since the density is uniform and purely advected, we expect the density solution at the steady state to match, up to numerical precision, the unit density field employed as initial guess.
We apply the variable density HHO formulation to a $h$-refined mesh sequence composed of regular triangular elements and compute numerical solutions for several polynomial degrees $k=0,1,2,3,4$. Triangular elements meshes are generated by means of the Delaunay algorithm and, at each step of the mesh sequence, a four fold increase of the mesh cardinality is approximately targeted. Errors in $L^{2}$-norm and $h$-convergence rates are computed testing against the analytical solution and tabulated for the approximated velocity, velocity gradients and pressure fields, see table 1. A k+1k{+}1 rate of error reduction is observed for the pressure error in $L^{2}$-norm. Since the flow is diffusion dominated at this Reynolds number, the velocity and velocity gradient errors in the $L^{2}$-norm show rates of convergence of k+2k{+}2 and k+1k{+}1, respectively. The errors in $L^{2}$-norm for the velocity divergence, also reported table 1, show that, the incompressibility constraint is enforced up to machine precision. Moreover, as confirmed by the density errors in $L^{2}$-norm reported in table 1, we are able to assess that a unit density field is obtained at all the polynomial degrees and on all the computations grids, compatibly with the claim that density is purely advected.
$\operatorname{card}\mathcal{T}_{h}$ $\mathfrak{E}\,{u_{h}}$ rate $\mathfrak{E}\,{\nabla u_{h}}$ rate $\mathfrak{E}\,{p_{h}}$ rate $\mathfrak{E}\,{\rho_{h}}$ $\mathfrak{E}\,{\nabla{\cdot}u_{h}}$
@@
Table 1: Kovasnzay test case: spatial convergence rates for $k=0,1,2,3,4$ HHO formulations over a Delaunay type $h$-refined triangular elements mesh sequence.
@@ 4.2 Guermond and Quartapelle manufactured solution
The spatial convergence results are reported in Table 2 for $k=0,1,2$ HHO formulations considering five grids of a triangular mesh sequence with $N=4\times{2^{i}}$, $i=0,\dots,5$. Time integration is performed relying on a third order accurate six stages ESDIRK scheme and using 320 uniform time steps. Note that, since the analytical pressure has a trigonometric spatial behavior, $k+1$ convergence rates for the pressure error in $L^{2}$-norm are attained at all polynomial degrees. Interestingly, the velocity and density errors in $L^{2}$-norm tabulated for $k=2$ are grid independent and allow to appreciate the temporal accuracy of the scheme. Indeed, since the by product between the density and the velocity is bilinear is $\bm{x}$, $k=2$ computations provide machine precision spatial accuracy for both the density and the velocity variables. This observation confirms pressure-robustness of the formulation, note that the velocity error is clearly independent from the pressure error.
The numerical results for $k=1$ show superconvergence of the velocity and density errors, in particular fourth order rate of error reduction is observed until spatial accuracy is on pair with temporal accuracy, resulting in error stagnation. The velocity error is slightly suboptimal at $k=0$. Nevertheless, these latter observations regarding spatial convergence rates must be carefully evaluated taking into account the very peculiar nature of the manufactured analytical solution.
$\mathfrak{E}\,{u_{h}}$ rate $\mathfrak{E}\,{\nabla u_{h}}$ rate $\mathfrak{E}\,{\rho_{h}}$ rate $\mathfrak{E}\,{p_{h}}$ rate $\mathfrak{E}\,{\nabla{\cdot}u_{h}}$
$k=0$,   $\operatorname{card}(\mathcal{T}_{h})=\left\{32,128,512,2048,8192,32768\right\}$
0.00398 0.0141 0.0917 0.0266 5.5e-16
0.00165 1.27 0.00580 1.28 0.0473 0.95 0.0146 0.86 4.7e-16
0.000774 1.09 0.00266 1.12 0.0244 0.96 0.00768 0.93 4.8e-16
0.000379 1.03 0.00129 1.05 0.0124 0.97 0.00393 0.97 4.9e-16
0.000188 1.01 0.000636 1.02 0.00633 0.98 0.00198 0.98 4.8e-16
6.31e-05 1.58 0.000190 1.74 0.00223 1.50 0.00106 0.90 4.9e-16
$k=1$,   $\operatorname{card}(\mathcal{T}_{h})=\left\{32,128,512,2048,8192,32768\right\}$
1.94e-06 7.08e-05 3.89e-07 0.00178 4.6e-16
1.23e-07 3.97 8.95e-06 2.98 1.72e-08 4.50 0.000443 2.00 5.6e-16
7.81e-09 3.99 1.12e-06 3.00 1.11e-09 3.95 0.000111 2.00 5.5e-16
5.19e-10 3.91 1.40e-07 3.00 2.53e-10 2.13 2.77e-05 2.00 5.5e-16
1.88e-10 1.46 1.78e-08 2.98 2.51e-10 0.01 6.93e-06 2.00 5.6e-16
1.86e-10 0.01 7.82e-09 1.19 2.71e-10 -0.11 1.20e-06 2.53 5.7e-16
$k=2$,   $\operatorname{card}(\mathcal{T}_{h})=\left\{32,128,512,2048,8192,32768\right\}$
2.04e-10 2.50e-09 1.79e-10 2.81e-05 6.9e-16
1.89e-10 3.40e-09 2.05e-10 3.57e-06 2.98 7.0e-16
1.87e-10 3.18e-09 2.33e-10 4.48e-07 2.99 7.0e-16
1.86e-10 3.10e-09 2.53e-10 5.63e-08 2.99 7.2e-16
1.86e-10 3.10e-09 2.69e-10 8.11e-09 2.79 7.0e-16
1.86e-10 3.10e-09 3.07e-10 4.42e-09 0.88 6.9e-16
@@
Table 2: Guermond and Quartapelle unsteady test case: spatial convergence rates for $k=0,1,2$ HHO formulations over a $h$-refined triangular elements mesh sequence. Time integration is conducted with a high-order accurate ESDIRK scheme, the time step is selected such that spatial errors dominate.
@@
Temporal convergence results are reported in Table 3 considering two ESDIRK schemes: i) the second order accurate three stages, so-called TR-BDF2, scheme, see [23] for details, ii) the fourth order accurate six stages scheme proposed in [22]. Both ESDIRK schemes are considered in combination with $k=2$ and $k=6$ HHO formulations on a triangular elements mesh with $N=8$ and $\operatorname{card}(\mathcal{T}_{h})=128$, the former provides machine precision spatial accuracy for the velocity and density variables while the latter guarantees negligible spatial errors for all variables, including the pressure. Note that, when considering the $k=2$ spatial discretization with third order accurate ESDIRK, the pressure error stagnates while, thanks to pressure-robustness, the velocity and density errors are almost identical to their $k=6$ counterparts.
@@
3 stages ESDIRK
k $\frac{7/4\pi}{\delta t}$ $\mathfrak{E}\,{u_{h}}$ rate $\mathfrak{E}\,{\nabla u_{h}}$ rate $\mathfrak{E}\,{\rho_{h}}$ rate $\mathfrak{E}\,{p_{h}}$ rate
6 20 0.000614 0.00161 0.00202 0.00168
40 0.000147 2.06 0.000389 2.06 0.000509 1.99 0.000404 2.07
80 3.62e-05 2.03 9.53e-05 2.03 0.000127 2.00 9.74e-05 2.04
160 8.98e-06 2.01 2.36e-05 2.01 3.19e-05 1.99 2.39e-05 2.02
320 2.23e-06 2.01 5.87e-06 2.01 7.98e-06 2.00 5.93e-06 2.01
2 20 0.000614 0.00161 0.00201 0.00168
40 0.000147 2.06 0.000388 2.06 0.000501 2.01 0.000401 2.07
80 3.62e-05 2.03 9.53e-05 2.03 0.000124 2.01 9.74e-05 2.04
160 8.98e-06 2.01 2.36e-05 2.01 3.10e-05 2.00 2.42e-05 2.01
320 2.23e-06 2.01 5.87e-06 2.01 7.73e-06 2.00 6.92e-06 1.81
6 stages ESDIRK
k $\frac{7/4\pi}{\delta t}$ $\mathfrak{E}\,{u_{h}}$ rate $\mathfrak{E}\,{\nabla u_{h}}$ rate $\mathfrak{E}\,{p_{h}}$ rate $\mathfrak{E}\,{\rho_{h}}$ rate
6 20 3.50e-06 1.26e-05 1.14e-05 1.03e-05
40 2.83e-07 3.63 1.43e-06 3.14 7.87e-07 3.87 1.22e-06 3.08
80 2.34e-08 3.60 1.74e-07 3.04 5.15e-08 3.93 1.66e-07 2.88
160 2.03e-09 3.53 2.26e-08 2.94 3.51e-09 3.88 2.50e-08 2.74
320 1.86e-10 3.45 3.11e-09 2.86 2.57e-10 3.77 4.06e-09 2.62
2 20 3.51e-06 1.26e-05 1.01e-05 1.10e-05
40 2.83e-07 3.63 1.43e-06 3.14 6.93e-07 3.87 3.78e-06 1.54
80 2.34e-08 3.60 1.77e-07 3.02 4.70e-08 3.88 3.58e-06 0.08
160 2.04e-09 3.52 2.40e-08 2.88 3.14e-09 3.90 3.57e-06 0.00
320 1.89e-10 3.43 3.40e-09 2.82 2.05e-10 3.93 3.57e-06 0.00
Table 3: Guermond and Quartapelle unsteady test case: temporal convergence rates for $k=2,6$ HHO formulations over a triangular elements mesh with $\operatorname{card}(\mathcal{T}_{h})=128$. Time integration is conducted with high-order accurate three and six stages ESDIRK schemes with increasingly small time steps, in particular $\delta t=\frac{7/4\pi}{N}$ with $N=20,40,80,160,320$.

### v3

@@ 4 Numerical results
In what follows we utilize the notation HHO-$\pi^{k}$ and HHO-$\pi^{k+1}$ to identify the two variants of the variable density HHO formulation introduced in see Sec. 3.4. In particular, HHO-$\pi^{k}$ and HHO-$\pi^{k+1}$ are based on the momentum equation spatial discretization $sd^{{\rm slm},\pi^{p}}$ with p=kp{=}k and p=k+1p{=}k{+}1, respectively. Moreover, for the sake of reporting convergence rates, we indicate with $\mathfrak{E}$ the error in $L^{2}$-norm computed against the analytical solution. Let $\bm{u}_{h}\in\mathcal{P}^{k+1}(\mathcal{T}_{h})\subset L^{2}(\Omega)$ be the discrete velocity obtained gluing together the cell components, the velocity error $\mathfrak{E}\;\bm{u}_{h}$ reads
$\mathfrak{E}\;\bm{u}_{h}\coloneqq\|\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)^{d}}=\left(\sum_{T\in\mathcal{T}_{h}}\int_{T}\left(\bm{u}-\bm{u}_{T}\right)\cdot\left(\bm{u}-\bm{u}_{T}\right)\right)^{\frac{1}{2}},$
and the velocity gradient error $\mathfrak{E}\;\nabla\bm{u}_{h}$ is computed as follows
$\mathfrak{E}\;\nabla\bm{u}_{h}\coloneqq\|\nabla\bm{u}-\nabla\bm{u}_{h}\|_{L^{2}(\Omega)^{d\times d}}=\left(\sum_{T\in\mathcal{T}_{h}}\int_{T}\left(\nabla\bm{u}-\nabla\bm{u}_{T}\right):\left(\nabla\bm{u}-\nabla\bm{u}_{T}\right)\right)^{\frac{1}{2}}.$
The pressure error $\mathfrak{E}\;p_{h}$, the density error $\mathfrak{E}\;\rho_{h}$ and the divergence error $\mathfrak{E}\;\nabla\cdot\bm{u}_{h}$ are computed in a similar fashion.
@@ 4.1 Kovasznay test case
$\operatorname{card}\mathcal{T}_{h}$ $\mathfrak{E}\,{\bm{u}_{h}}$ rate $\mathfrak{E}\,{\nabla\bm{u}_{h}}$ rate $\mathfrak{E}\,{p_{h}}$ rate $\mathfrak{E}\,{\rho_{h}}$ $\mathfrak{E}\,{\nabla{\cdot}\bm{u}_{h}}$
$k=0$
50 0.677 5.818 0.666103 9.8e-16 1.8-16
192 0.372 0.89 4.158 0.50 0.330301 1.04 8.5e-16 2.5-16
810 0.188 0.95 2.267 0.84 0.153069 1.07 1.4e-15 2.7-16
3126 0.100 0.93 1.244 0.89 0.0757808 1.04 1.3e-15 2.6-16
12706 0.0504 0.98 0.629 0.97 0.0354129 1.09 1.6e-15 2.6-16
$k=1$
50 0.198 2.986 0.111 8.2e-14 3.4e-16
192 0.0483 2.10 1.160 1.41 0.0238 2.29 1.5e-14 3.6e-16
810 0.00793 2.51 0.380 1.55 0.00440 2.35 4.5e-14 3.2e-16
3126 0.00106 2.98 0.112 1.80 0.000881 2.38 7.0e-14 3.3e-16
12706 0.000142 2.86 0.0307 1.85 0.000202 2.10 9.1e-14 3.2e-16
$k=2$
50 0.0418 1.041 0.0286 2.3-14 4.3e-16
192 0.00417 3.43 0.185 2.56 0.00227 3.77 5.3-14 3.8e-16
810 0.000303 3.64 0.0276 2.65 0.000202 3.36 4.3-14 3.8e-16
3126 2.53e-05 3.67 0.00460 2.65 2.53e-05 3.07 3.6-14 3.8e-16
12706 1.64e-06 3.90 0.000589 2.93 3.23e-06 2.94 1.3-13 3.8e-16
$k=3$
50 0.00827 0.266 0.00198 5.4e-14 4.0e-16
192 0.000383 4.57 0.0261 3.45 0.000150 3.83 2.9e-14 4.2e-16
810 2.02e-05 4.09 0.00243 3.30 1.23e-05 3.47 2.3e-14 3.6e-16
3126 6.35e-07 5.13 0.000163 4.00 8.05e-07 4.05 5.9e-14 3.9e-16
12706 2.29e-08 4.74 1.17e-05 3.76 5.13e-08 3.93 7.9e-14 3.8e-16
$k=4$
50 0.000933 0.0453 0.000487 2.4e-14 4.8e-16
192 2.31e-05 5.49 0.00227 4.45 1.19e-05 5.52 4.2e-14 4.8e-16
810 5.19e-07 5.28 9.29e-05 4.45 4.28e-07 4.62 5.3e-14 4.3e-16
3126 1.17e-08 5.62 4.15e-06 4.60 1.47e-08 4.99 1.0e-13 4.5e-16
12706 1.98e-10 5.81 1.37e-07 4.86 5.14e-10 4.79 5.3e-13 4.4e-16
Table 1: Kovasnzay test case: spatial convergence rates for $k=0,1,2,3,4$ HHO-$\pi^{k}$ formulations over a Delaunay type $h$-refined triangular elements mesh sequence.
$\operatorname{card}\mathcal{T}_{h}$ $\mathfrak{E}\,{\bm{u}_{h}}$ rate $\mathfrak{E}\,{\nabla\bm{u}_{h}}$ rate $\mathfrak{E}\,{p_{h}}$ rate $\mathfrak{E}\,{\rho_{h}}$ $\mathfrak{E}\,{\nabla{\cdot}\bm{u}_{h}}$
@@
Table 2: Kovasnzay test case: spatial convergence rates for $k=0,1,2,3,4$ HHO-$\pi^{k+1}$ formulations over a Delaunay type $h$-refined triangular elements mesh sequence.
As a first validation step we seek to numerically demonstrate pure advection of the density by tackling the well-known Kovasznay flow problem [19]. The analytical solution of the incompressible Navier–Stoke equations, which describes steady flow behind a grid made of equally spaces parallel rods for a fluid with constant density, reads as follows
$\bm{u}=\left[1-\mathrm{e}^{\kappa x}\cos{\left(2\pi y\right)}\right]\bm{i}+\dfrac{\kappa}{2\pi}\mathrm{e}^{\kappa x}\sin{\left(2\pi y\right)}\bm{j},\qquad p=p_{0}-\dfrac{1}{2}\mathrm{e}^{2\kappa x},\qquad\rho=1,$ (32)
$\kappa=\dfrac{\mathrm{Re}}{2}-\sqrt{\dfrac{\mathrm{Re}^{2}}{4}+4\pi^{2}}.$
We define the Kovasznay flow problem on the bi-unit square computational domain $\Omega=\left(-0.5,1.5\right){\times}\left(0,2\right)$ and derive boundary conditions from the analytical solution (32) at $\mathrm{Re}{=}\frac{1}{\mu}{=}40$. In particular, a Neumann boundary condition is imposed at the outflow boundary (right side of the square domain) while Dirichlet boundary conditions are imposed on the remaining boundaries. We apply the variable density HHO-$\pi^{k}$ and HHO-$\pi^{k+1}$ formulations to a $h$-refined mesh sequence composed of regular triangular elements and compute numerical solutions for several polynomial degrees $k=0,1,2,3,4$. Triangular elements meshes are generated by means of the Delaunay algorithm and, at each step of the mesh sequence, a four fold increase of the mesh cardinality is approximately targeted.
Tables 1 and 2 report the results obtained by means of HHO-$\pi^{k}$ and HHO-$\pi^{k+1}$ formulations, respectively. Errors in $L^{2}$-norm and $h$-convergence rates are computed testing against the analytical solution and tabulated for the approximated velocity, velocity gradients and pressure fields. A k+1k{+}1 rate of error reduction is observed for the pressure error in $L^{2}$-norm. Since the flow field is diffusion dominated at this Reynolds number, the velocity and velocity gradient errors in the $L^{2}$-norm are expected to exhibit convergence rates of k+2k{+}2 and k+1k{+}1, respectively. The major difference between HHO-$\pi^{k+1}$ and HHO-$\pi^{k}$ is observed at $k=0$, note that both the velocity and the velocity gradient are first order accurate in case of HHO-$\pi^{k}$ while HHO-$\pi^{k+1}$ is able to reach second order accuracy for the velocity unknown. When considering $k>0$, both HHO-$\pi^{k+1}$ and HHO-$\pi^{k}$ are able to reach the theoretical convergence rates. Nonetheless, HHO-$\pi^{k+1}$ is the best accurate scheme and the asymptotic convergence rates are usually attained on coarser meshes.
The errors in $L^{2}$-norm for the velocity divergence and the density, reported in Tables 1-2, show that, since the incompressibility constraint is enforced up to machine precision, the velocity field is pointwise divergence-free and, since the density is purely advected, a unit density field is obtained irrespectively from the discretization parameters. As expected, the density solution at the steady state matches, up to numerical precision, the unit density field employed as initial guess.
@@ 4.2 Guermond and Quartapelle manufactured solution
$\operatorname{card}\mathcal{T}_{h}$ $\mathfrak{E}\,{\bm{u}_{h}}$ rate $\mathfrak{E}\,{\nabla\bm{u}_{h}}$ rate $\mathfrak{E}\,{\rho_{h}}$ rate $\mathfrak{E}\,{p_{h}}$ rate $\mathfrak{E}\,{\nabla{\cdot}\bm{u}_{h}}$
$k=0$
32 0.00773 0.0297 0.0628 0.0327 5.4e-16
128 0.00397 0.96 0.0151 0.97 0.0328 0.93 0.0168 0.96 4.9e-16
512 0.00201 0.98 0.00760 1.00 0.0172 0.93 0.00870 0.95 4.6e-16
2048 0.001017 0.99 0.00379 1.00 0.00897 0.94 0.00444 0.97 4.9e-16
8192 0.000511 0.99 0.00189 1.00 0.00462 0.96 0.00225 0.98 4.9e-16
32768 0.000256 1.00 0.000943 1.00 0.00236 0.97 0.00113 0.99 4.8e-16
131072 0.000128 1.00 0.000471 1.00 0.00120 0.98 0.000567 1.00 4.8e-16
$k=1$
32 3.79e-06 0.000124 8.69e-07 0.00122 7.2e-16
128 2.49e-07 3.93 1.62e-05 2.94 6.37e-08 3.77 0.000307 1.99 7.3e-16
512 1.59e-08 3.97 2.07e-06 2.97 4.38e-09 3.86 7.69e-05 2.00 6.9e-16
2048 1.01e-09 3.97 2.61e-07 2.99 4.12e-10 3.41 1.92e-05 2.00 6.8e-16
8192 1.94e-10 2.39 3.29e-08 2.99 2.55e-10 0.69 4.81e-06 2.00 6.9e-16
32768 1.87e-10 0.05 5.15e-09 2.68 2.71e-10 -0.09 1.20e-06 2.00 6.8e-16
$k=2$
32 2.06e-10 2.78e-09 1.77e-10 5.65e-05 6.8e-16
128 1.90e-10 3.66e-09 2.05e-10 7.07e-06 3.00 6.8e-16
512 1.87e-10 3.19e-09 2.32e-10 8.83e-07 3.00 6.8e-16
2048 1.86e-10 3.10e-09 2.53e-10 1.10e-07 3.00 6.8e-16
8192 1.86e-10 3.10e-09 2.70e-10 1.44e-08 2.94 6.9e-16
32768 1.86e-10 3.10e-09 3.07e-10 4.42e-09 1.70 6.8e-16
@@
Table 3: Guermond and Quartapelle unsteady test case: spatial convergence rates for $k=0,1,2$ HHO-$\pi^{k}$ formulations over a $h$-refined triangular elements mesh sequence. Time integration is conducted with a high-order accurate ESDIRK scheme, the time step is selected such that spatial errors dominate.
$\operatorname{card}\mathcal{T}_{h}$ $\mathfrak{E}\,{\bm{u}_{h}}$ rate $\mathfrak{E}\,{\nabla\bm{u}_{h}}$ rate $\mathfrak{E}\,{\rho_{h}}$ rate $\mathfrak{E}\,{p_{h}}$ rate $\mathfrak{E}\,{\nabla{\cdot}\bm{u}_{h}}$
$k=0$
32 0.00249 0.00721 0.0609 0.0313 5.1e-16
128 0.00112 1.15 0.00318 1.18 0.0317 0.94 0.0158 0.98 4.8e-16
512 0.000533 1.07 0.00152 1.06 0.0165 0.94 0.00801 0.98 5.1e-16
2048 0.000259 1.04 0.000752 1.02 0.00857 0.95 0.00404 0.99 4.8e-16
8192 0.000128 1.02 0.000376 1.00 0.00440 0.96 0.00203 0.99 4.9e-16
32768 6.38e-05 1.01 0.000189 0.99 0.00224 0.97 0.00102 1.00 4.8e-16
131072 3.19e-05 1.00 9.51e-05 0.99 0.00114 0.98 0.000509 1.00 4.9e-16
$k=1$
32 3.88e-06 0.000126 8.70e-07 0.00122 6.6e-16
128 2.52e-07 3.94 1.64e-05 2.95 6.37e-08 3.77 0.000307 1.99 6.6e-16
512 1.60e-08 3.98 2.08e-06 2.98 4.37e-09 3.86 7.69e-05 2.00 7.0e-16
2048 1.02e-09 3.98 2.62e-07 2.99 4.12e-10 3.41 1.92e-05 2.00 6.8e-16
8192 1.94e-10 2.39 3.30e-08 2.99 2.56e-10 0.69 4.81e-06 2.00 6.8e-16
32768 1.87e-10 0.05 5.15e-09 2.68 2.71e-10 -0.09 1.20e-06 2.00 6.8e-16
$k=2$
32 2.06e-10 2.78e-09 1.77e-10 5.65e-05 7.0e-16
128 1.90e-10 3.67e-09 2.05e-10 7.07e-06 3.00 6.9e-16
512 1.87e-10 3.19e-09 2.32e-10 8.84e-07 3.00 7.1e-16
2048 1.86e-10 3.10e-09 2.53e-10 1.10e-07 3.00 6.9e-16
8192 1.86e-10 3.10e-09 2.70e-10 1.44e-08 2.94 6.8e-16
32768 1.86e-10 3.10e-09 3.07e-10 4.42e-09 1.70 6.9e-16
Table 4: Guermond and Quartapelle unsteady test case: spatial convergence rates for $k=0,1,2$ HHO-$\pi^{k+1}$ formulations over a $h$-refined triangular elements mesh sequence. Time integration is conducted with a high-order accurate ESDIRK scheme, the time step is selected such that spatial errors dominate.
The spatial convergence results are reported in Table 3 and Table 4 for $k=0,1,2$ HHO-$\pi^{k}$ and HHO-$\pi^{k+1}$ formulations, respectively, considering five grids of a triangular mesh sequence with $N=4\times{2^{i}}$, $i=0,\dots,5$. Time integration is performed relying on a third order accurate six stages ESDIRK scheme and using 320 uniform time steps. Note that, since the analytical pressure has a trigonometric spatial behavior, $k+1$ convergence rates for the pressure error in $L^{2}$-norm are attained at all polynomial degrees. Interestingly, the velocity and density errors in $L^{2}$-norm tabulated for $k=2$ are grid independent and allow to evaluate the temporal accuracy of the scheme. Indeed, since the by product between the density and the velocity is a second degree polynomial in $\bm{x}$, $k=2$ computations provide machine precision spatial accuracy for both the density and the velocity variables. This behavior confirms that the HHO formulations are pressure-robust, note indeed that the velocity error is clearly independent from the pressure error.
Overall HHO-$\pi^{k+1}$ and HHO-$\pi^{k}$ formulations provide almost identical accuracy and convergence rates for $k=1,2$. The numerical results for $k=1$ show fourth order rate of error reduction for both the velocity and density unknowns, until spatial accuracy is on pair with temporal accuracy, resulting in error stagnation. We remark that, spatial convergence rates for $k=1$ must be carefully evaluated considering that, according to the manufactured solution, both the velocity and the density unknowns are linear functions of $\bm{x}$. At $k=0$, both the velocity and the velocity gradient exhibiting first order of convergence in $L^{2}$-norm, and, similarly to the Kovasznay test case, HHO-$\pi^{k+1}$ shows improved accuracy with respect to HHO-$\pi^{k}$.
3 stages ESDIRK
k $\frac{7/4\pi}{\delta t}$ $\mathfrak{E}\,{\bm{u}_{h}}$ rate $\mathfrak{E}\,{\nabla\bm{u}_{h}}$ rate $\mathfrak{E}\,{\rho_{h}}$ rate $\mathfrak{E}\,{p_{h}}$ rate $\mathfrak{E}\,{\nabla{\cdot}\bm{u}_{h}}$
6 20 0.000614 0.00161 0.00202 0.00168 8.7e-17
40 0.000147 2.06 0.000389 2.06 0.000509 1.99 0.000404 2.07 8.2e-17
80 3.62e-05 2.03 9.53e-05 2.03 0.000127 2.00 9.74e-05 2.04 8.6e-17
160 8.98e-06 2.01 2.36e-05 2.01 3.19e-05 1.99 2.39e-05 2.02 8.9e-17
320 2.23e-06 2.01 5.87e-06 2.01 7.98e-06 2.00 5.93e-06 2.01 8.8e-17
2 20 0.000614 0.00161 0.00201 0.00168 6.6e-17
40 0.000147 2.06 0.000388 2.06 0.000501 2.01 0.000401 2.07 6.4e-17
80 3.62e-05 2.03 9.53e-05 2.03 0.000124 2.01 9.74e-05 2.04 7.2e-17
160 8.98e-06 2.01 2.36e-05 2.01 3.10e-05 2.00 2.42e-05 2.01 7.3e-17
320 2.23e-06 2.01 5.87e-06 2.01 7.73e-06 2.00 6.92e-06 1.81 6.3e-17
6 stages ESDIRK
k $\frac{7/4\pi}{\delta t}$ $\mathfrak{E}\,{\bm{u}_{h}}$ rate $\mathfrak{E}\,{\nabla\bm{u}_{h}}$ rate $\mathfrak{E}\,{p_{h}}$ rate $\mathfrak{E}\,{\rho_{h}}$ rate $\mathfrak{E}\,{\nabla{\cdot}\bm{u}_{h}}$
6 20 3.50e-06 1.26e-05 1.14e-05 1.03e-05 8.1e-16
40 2.83e-07 3.63 1.43e-06 3.14 7.87e-07 3.87 1.22e-06 3.08 8.4e-16
80 2.34e-08 3.60 1.74e-07 3.04 5.15e-08 3.93 1.66e-07 2.88 8.4e-16
160 2.03e-09 3.53 2.26e-08 2.94 3.51e-09 3.88 2.50e-08 2.74 8.2e-16
320 1.86e-10 3.45 3.11e-09 2.86 2.57e-10 3.77 4.06e-09 2.62 7.4e-16
2 20 3.51e-06 1.26e-05 1.01e-05 1.10e-05 7.1e-16
40 2.83e-07 3.63 1.43e-06 3.14 6.93e-07 3.87 3.78e-06 1.54 7.0e-16
80 2.34e-08 3.60 1.77e-07 3.02 4.70e-08 3.88 3.58e-06 0.08 7.3e-16
160 2.04e-09 3.52 2.40e-08 2.88 3.14e-09 3.90 3.57e-06 0.00 7.3e-16
320 1.89e-10 3.43 3.40e-09 2.82 2.05e-10 3.93 3.57e-06 0.00 7.2e-16
Table 5: Guermond and Quartapelle unsteady test case: temporal convergence rates for $k=2,6$ HHO-$\pi^{k+1}$ formulations over a triangular elements mesh with $\operatorname{card}(\mathcal{T}_{h})=128$. Time integration is conducted with high-order accurate three and six stages ESDIRK schemes with increasingly small time steps, in particular $\delta t=\frac{7/4\pi}{N}$ with $N=20,40,80,160,320$.
@@
Temporal convergence results are reported in Table 5 considering HHO-$\pi^{k+1}$ (HHO-$\pi^{k}$ is omitted as it provides similar results) and two ESDIRK schemes: i) the second order accurate three stages, so-called TR-BDF2, scheme, see [25] for details, ii) the fourth order accurate six stages scheme proposed in [24]. Both ESDIRK schemes are considered in combination with $k=2$ and $k=6$ formulations on a triangular elements mesh with $N=8$ and $\operatorname{card}(\mathcal{T}_{h})=128$, the former provides machine precision spatial accuracy for the velocity and density variables while the latter guarantees negligible spatial errors for all variables, including the pressure. Note that, when considering the $k=2$ spatial discretization with third order accurate ESDIRK, the pressure error stagnates while, thanks to pressure-robustness, the velocity and density errors are almost identical to their $k=6$ counterparts.